OpenStudy (anonymous):

multiply and simplify by factoring sqrt45a^9b x sqrt27a^10b^8

OpenStudy (asnaseer):

is this your equation?\[\sqrt{45a^9b}*\sqrt{27a^{10}b^8}\]

OpenStudy (anonymous):

yes :D

OpenStudy (asnaseer):

first note that:\[\sqrt{x}*\sqrt{y}=\sqrt{xy}\]so we get:\[\sqrt{45a^9b}*\sqrt{27a^{10}b^8}=\sqrt{45*27*a^9*a^{10}*b*b^8}\]\[=\sqrt{3^2*5*3^3*a^{19}*b^9}\]\[=\sqrt{3^5*5*a^{19}b^9}\]do you know how to simplify from here on?

OpenStudy (anonymous):

no clue :(

OpenStudy (asnaseer):

ok, what do you think the answer to this is? \[\sqrt{x^{2n}}\]

OpenStudy (anonymous):

uhm... x times x=2xn?

OpenStudy (asnaseer):

ok - you really need to learn the rules of indexes first, try here: http://www.mathsteacher.com.au/year8/ch07_indices/02_mult/law.htm

OpenStudy (asnaseer):

another useful resource: http://www.revisesmart.co.uk/maths/core-1/indices-and-surds.html

jhonyy9 (jhonyy9):

this will be 3a4sqrt5ab * 3a5b4sqrt3 = 9a9b4sqrt15ab

jhonyy9 (jhonyy9):

the second 9 and 4 are exponents

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

\[9 \sqrt{15} a^{19/2} b^{9/2} \]